OpenStudy (anonymous):

-3x + 4/7y = 20 y = 14 answers (14, -4) (-4, 14) (14, -4) (-4, 14)

OpenStudy (anonymous):

Please help fast

OpenStudy (anonymous):

please help link,zoro:

OpenStudy (anonymous):

I think its the second one

OpenStudy (anonymous):

im most definetly giving you a medal

OpenStudy (anonymous):

what is the kind of math your doing is it equations that need to be graph

OpenStudy (anonymous):

solving the system of equations.

OpenStudy (anonymous):

its a pretest but I need to make a good grade better yet I have to

OpenStudy (anonymous):

im homeschooled

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so get this they wanted me to do this survey for my school and its working in module 7 and im in 3 that's why I don't get this stupid question

OpenStudy (anonymous):

if y =14 4/7*14 =8 8+(-3) =5 5+15 =20 I think its the first one

OpenStudy (anonymous):

yes it is thanks